If the price level increases by 0.2 percent for every $100 billion increase in the money supply, by how much might prices rise if the Fed increases total reserves by $80 billion and the reserve requirement is 0.05?
Answer #1

Answer

Money multiplier(m) = 1/rr where rr = required reserve ratio = 0.05 => m = 1/0.05 = 20.

Interpretation of m :

$1 increase in reserves will result in increase by $20. Thus 80 billion increase in reserves will increase money supply(M) by 20*80 billion = 1600 billion.

Here it is given that $100 billion will increase Price level by 0.2% , thus $1600 billion will increase Price level by 0.2*16 = 3.2%

Hence, Price level will rise by 3.2%.